Q1: Using Simpsons 1/3 rule to solve Lo®3x?dx
Answer #1

I have completed this problem Please give thumbs up if you like it

Stept: Simpsons Rule formula for n=2 en ДХ 3 [FC%) +45C%) +2f6x)] Ax=b-a 3-0 = 2 3) are The value of the quadisatic function